National Income Accounts.
X 100. Transfers 100. PIT 220. SST 60.
IBT1 40. NPI 40. I 100. CRE 20. C 700.
NFFI 100. Depreciation 60. IBT2 40. G 300. M 200.
Determine GNP.
1
Expert's answer
2016-05-17T12:39:03-0400
The gross national product is defined as follows:
GNP=GDP+NFFI
GDP=C+I+G+(X-M)
GDP = 700 + 100 + 300 + (100-200) =1000
GNP=1000+100=1100